OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). cos^2x + 2 cos x + 1 = 0

OpenStudy (anonymous):

Factorise: \[\cos^2x+2\cos x+1=0\\ (\cos x+1)^2=0\] So you have \[\cos x=-1\] Now, which value of \(x\) in \([0,2\pi)\) gives you \(\cos x=-1\)?

OpenStudy (anonymous):

pi

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

so that it?

OpenStudy (anonymous):

Yes, unless you can think of any other \(x\) that make \(\cos x=-1\) ?

OpenStudy (anonymous):

alright thanks

OpenStudy (anonymous):

you're welcome